OpenStudy (anonymous):

The area of a garden is given by the trinomial g2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width

OpenStudy (anonymous):

g - 6

OpenStudy (anonymous):

You have g^2-2g-24=(g+4)*width So the width has to be g+or- a number to get something starting with g^2 It has -24 so the number has to be -6 as 4*-6=-24 So the answer is g-6 You can check by expanding the bracket (g+4)(g-6)

jhonyy9 (jhonyy9):

g^2 -2g -24=(g-6)(g+4)

jhonyy9 (jhonyy9):

hello andrás

OpenStudy (anonymous):

hi :-)

jhonyy9 (jhonyy9):

pardon András !!!

jhonyy9 (jhonyy9):

happy new year azaz Boldog Új Évet !!!